OpenStudy (anonymous):

What are the x and y intercepts of the quadraric equation y=(x-3)^2? identify the x intercepts by graphing the quadratric equation

OpenStudy (campbell_st):

the x intercept is when y = 0 (x - 3)^2 = 0 so solve x - 3 = 0 y- intercept x = 0 so y = (0 -3)^2 y = 9